OpenStudy (anonymous):

solve for x log5x=-3

OpenStudy (anonymous):

log(5x) = -3 like that?

OpenStudy (anonymous):

Cancel the logarithms by taking the exponential of both sides.

OpenStudy (anonymous):

\[5x= \frac{ 1 }{ e^3 } => x = \frac{ 1 }{ 5e^3 }\]

OpenStudy (anonymous):

the 5 is in between the log and th x but its small and on the bottom, i just didnt know how to type it

OpenStudy (anonymous):

That made no sense, but good luck, you should know what to do with the above example.

OpenStudy (anonymous):

5x=-3 x=-3/5 x=-.6